OpenStudy (anonymous):

Evaluate: integral of [(1)/sqrt(-x^2-4x-3)]dx

OpenStudy (anonymous):

dat is hard

OpenStudy (anonymous):

can u tell me a polyhedron in an easy way

OpenStudy (sirm3d):

\[\Large \sqrt{-x^2-4x-3}=\sqrt{1-(x+2)^2}\] you may use trigonometric substitution if you haven't recognize the integral as Arcsine or \(sin^{-1}\)

OpenStudy (anonymous):

ohhhhhh!

OpenStudy (anonymous):

xD wow. now i see it

OpenStudy (anonymous):

so the answer would just be arcsin (x+2) ?

OpenStudy (sirm3d):

yup.